OpenStudy (anonymous):

Find the Product of the following matrices, if possible: will give a medal :)

OpenStudy (anonymous):



OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

I could do this but I'd have to do it by hand and you might not follow it. This video might be more useful, it's multiplying a 3x3 matrix by a 3x1 just as you are: http://www.youtube.com/watch?v=IRdVU8tyRDU

OpenStudy (anonymous):

\[\left[\begin{matrix}4&1&0\\-1&1&-4\\-8&9&-9\end{matrix}\right]\cdot\left[\begin{matrix}6\\1\\8\end{matrix}\right]=\left[\begin{matrix}(4\times6)+(1\times1)+(0\times8)\\ (-1\times6)+(1\times1)+(-4\times8)\\(-8\times6)+(9\times1)+(-9\times8)\end{matrix}\right]\]

OpenStudy (anonymous):

thank you so much @SithsAndGiggles

OpenStudy (anonymous):

thank you so much @SithsAndGiggles

OpenStudy (anonymous):

You're welcome.